# Ultraviolet spectroscopy reveals a hot and luminous companion to the Be star + black hole candidate MWC 656

### v1

@@ 5 Signs of the companion in the UV
The HST/STIS FUV spectra, obtained at four orbital phases (two near opposite quadratures at , and two near conjunction, when the Be star was in front of the companion along the line of sight at , see Fig. 2), are displayed in Fig. 4. The Be star model reproduces most of the FUV spectrum well, confirming that the observed flux is dominated by the Be star. However, several features are not reproduced by the model and instead suggest the presence of a hotter component in the system. We discuss these features in the following.
@@ 5.1 He ii emission
The He ii line, shown in the rightmost zoom panel of Fig. 4, displays clear phase-dependent variability (see also Fig. 11 in the appendix). Mid- to late-type B stars generally do not exhibit He ii lines, and therefore the variability seen here is unlikely to originate from the Be star. Despite the limited number of available epochs, we detect an emission component that shifts in anti-phase with the Be star, consistent with the companion (see Appendix B). This behaviour mirrors our findings in the optical, where the He ii line also appears in emission, tracking the companion.
@@ 5.2 High-ionisation wind lines
Beyond He ii, the spectra show several high-ionisation species, shown in Fig. 4. The N v 1239, 1243 doublet exhibits prominent P Cygni profiles, especially near quadrature, with blue-shifted absorption and red-shifted emission components. This is suggestive of a hot, radiatively driven wind. Although Si iv 1393, 1402 and C iv 1548, 1551 lines do not show clear emission components, they exhibit sloped blue-shifted absorption troughs indicative of wind outflows, especially at and . These wind features coincide with an interstellar absorption component.
The Si iv and C iv lines can arise from superionisation in B-star winds, typically interpreted as the result of shock-generated X-rays that raise ionisation levels above those expected for the stellar temperature (e.g. Puls et al., 2008). Such effects are common in Be stars and explain their stronger UV wind lines compared to non-emission B stars (Grady et al., 1987; Prinja, 1989). Indeed, incorporating the observed level of X-rays into our 21 kK Be-star model reproduces the absorption components of the Si iv and C iv resonance lines. However, even accounting for superionisation, N v P-Cygni emission is only observed in the most luminous Be stars earlier than B1 (Kogure and Hirata, 1982; Rivinius et al., 2013). As discussed in Sect. 8.1, no comparable N v emission is seen in other Be stars of similar type to MWC 656. Producing such features in MWC 656 therefore requires either a substantially hotter ionising source or an unusually high X-ray luminosity.
@@ 5.3 Doppler shifts
Using the orbital solution presented in Sect. 3.2, we predicted Doppler shifts for both components across the observations. For the Be star, the expected maximum RV amplitude between observations is approximately 30 km s-1. To estimate RV shifts from the observed UV spectra, we performed an iterative cross-correlation of each epoch against the spectrum, focussing on wavelength regions free from strong interstellar absorption and wind lines. This yields relative RVs of , , and km s-1 for , respectively. The observed shift at is qualitatively consistent with the expected lower RV at this phase, although the measured amplitude is smaller than predicted. However, we note that, given the moderate S/N, broad spectral features, and relatively low RV amplitude of the Be star, precise measurements are challenging. The companion is predicted to have an RV amplitude of about 140 km s-1 between observations. However, the FUV spectrum is dominated by the Be star, making it challenging to directly measure the companion’s RVs. Nevertheless, we recover He ii emission from the companion through spectral disentangling using the predicted orbital velocities (Fig. 11), providing indirect support for the expected motion. A puzzling feature of the UV observations is the apparent Doppler shift seen in the high-ionisation wind lines, which follows neither of the binary components and is difficult to explain in any binary companion scenario. We address this aspect in more detail in Sect. 8.2.
@@ 5.4 Phase-dependent spectral variability
The four epochs of HST/STIS FUV spectroscopy reveal phase-dependent variability in the high-ionisation UV wind lines. Resonance line profiles appear to strengthen near quadrature, with, for example, the C iv doublet showing more prominent blue-shifted absorption troughs during these phases, while features weaken near conjunction (–0.15). Variability in the line profiles of resonance lines such as Si iv and C iv is common in Be stars (e.g. Grady et al., 1987). But the observed UV line profile changes may also reflect the interaction between the companion and the Be disc. In binaries, hydrodynamic simulations show that the companion can perturb the Be disc, exciting spiral density waves that produce phase-dependent variations in emission and absorption (Panoglou et al., 2016; Rubio et al., 2025). The phase-dependence could also indicate ionisation of the Be star wind by the hot companion, in which case the absorption would originate on the far side of the system during conjunction, or the wind of the hot companion itself, whose signatures diminish when it is partially obscured behind the Be star. However, we note that with only four UV epochs, some apparent phase-dependence could be coincidental.
@@ 6 Composite model with a stripped-star companion
The UV data revealed spectral features, such as high ionisation wind lines in N v and He ii emission, absent in the Be star spectrum, which point toward the presence of a hotter luminous companion in the system (Sect. 5). Motivated by these findings, we modelled the binary as a composite of two luminous stars, featuring the Be star and a stripped helium star companion. We performed a combined spectroscopic and photometric analysis, with a setup similar to that in Sect. 4, with the companion star parameters optimised to reproduce the observed UV wind lines.
@@
The companion’s effective temperature is primarily constrained by the observed UV emission lines. The presence of N v requires a minimum temperature of kK. However, reproducing the observed N v emission without accompanying strong C iv or Si iv emission requires significantly higher temperatures, kK, consistent with trends seen in Wolf-Rayet stars (e.g., Willis et al., 1986; Hamann and Gräfener, 2004; Hamann et al., 2006; Sander et al., 2025).
We explored companion temperatures from 65 to 95 kK, with corresponding scaled between 4.6 and 5.6, to be consistent with the helium Zero-Age Main Sequence (He-ZAMS). Models with lower temperatures produce stronger C iv and Si iv lines but weaker N v features. In contrast, models that are too hot have a steep SED, which reduces their contribution to the optical spectrum. By balancing the strength and shape of the N v P Cygni profile with the contribution to optical lines such as He ii 4686, our models suggest a best-fit temperature of kK. The sensitivity of the fit to the assumed temperature is illustrated in Appendix E, where we compare cooler and hotter companion models to the observations.
The resulting best-fit composite model is shown in Fig. 6, with key parameters summarised in Table 2. We derive mag and . We note that the corresponding -band extinction of mag agrees well with 3D dust map predictions (e.g., from Bayestar19; Green, 2018; Green et al., 2019), suggesting negligible system-intrinsic extinction. The Be star has inferred luminosity and stellar radius , the latter determined using the Stefan-Boltzmann law. As expected, attributing part of the observed flux to the stripped companion results in slightly lower inferred luminosity and radius compared to the single-star model (Sect. 4), but the differences are minimal. This is because the companion’s contribution to the total flux is small, and the Be star dominates both observed UV and optical light.
@@ 7.1 Companion Mass
The mass of the companion in MWC 656 can be constrained by combining the dynamical mass ratio with spectroscopic and evolutionary estimates of the Be star mass.
@@
illustrated in Fig. 7. The Be star mass inferred from our composite spectroscopic model is (The mass obtained from the single-star analysis was 8 ). Combining this with the mass function implies an inclination of degrees and a companion mass of .
An evolutionary mass was also derived by comparing the effective temperature and luminosity of the Be star with simulated evolutionary tracks. We used the BONNSAI444The BONNSAI web-service is available at www.astro.uni-bonn.de/stars/bonnsai code (Schneider et al., 2014), which implements a Bayesian statistical method to determine probability distributions of fundamental stellar parameters. We used the massive star evolutionary model grids by Brott et al. (2011) and flat prior distributions in stellar mass and age. This yields , corresponding to a companion mass of – consistent with the spectroscopic mass estimate. However, this value may represent an upper limit, as past mass accretion could make the Be star temporarily appear over-luminous relative to its true mass (e.g. Kippenhahn and Meyer-Hofmeister, 1977; Lau et al., 2024).
Taken together, the dynamical, spectroscopic, and evolutionary constraints consistently place the companion mass between 1.1 and 1.9 M⊙. Two key conclusions follow: First, a BH companion is strongly disfavoured. The inferred mass lies well below the canonical lower mass threshold for BHs (3 M⊙). For comparison, C14 proposed a BH mass of 3.8–6.9 M⊙. Accommodating even the lowest of these values would require a Be star mass of 18.5 M⊙, which is ruled out at by both spectroscopic and evolutionary estimates. Second, a white dwarf companion is also unlikely. Even the minimum allowed dynamical companion mass at ( M⊙) implies a very massive white dwarf, and the most likely companion mass of M⊙ exceeds the Chandrasekhar limit (1.4 M⊙), placing most of the plausible mass range above this threshold.
@@ 7.2 Could the companion still be an accreting compact object?
A white dwarf accretor provides another possible explanation. X-rays from accretion could photoionise the Be-star wind and produce N v emission. This scenario has been proposed to explain the so-called Cas stars, Be stars with anomalously high X-ray luminosities that may host accreting white dwarfs (e.g. Smith et al., 2016; Gies et al., 2023). Although MWC 656 is significantly less X-ray luminous than typical Cas stars, we tested this scenario by incorporating increasing levels of X-ray heating in the Be-star model (Sect. 5.2). Reproducing the observed N v strength requires X-ray luminosities about two orders of magnitude higher than observed in MWC 656, and the models simultaneously predict strong C iv and Si iv emission that is not seen in the data.
@@ 8.1 Comparison of MWC 656 to known stripped stars
To place our results in context, we compare the inferred properties of the stripped star in MWC 656 to those of known stripped stars. These hot, compact helium stars are formed through envelope stripping in binary systems. Stripped stars span a wide mass range, from low-mass subdwarf B and O stars (; e.g. Heber, 2009; Mourard et al., 2015; Klement et al., 2022) to massive Wolf-Rayet stars (; e.g., Shenar et al., 2016, 2018, 2020b; Sander and Vink, 2020), with several intermediate-mass cases recently identified in the Magellanic Clouds (Ludwig et al., 2025; Drout et al., 2023; Ramachandran et al., 2023, 2024).
Stripped stars are often found with rapidly rotating main-sequence companions, believed to have spun up via accretion during envelope stripping. Wang et al. (2021), for instance, identified Galactic Be+subdwarf candidates using HST UV spectroscopy, detecting hot subdwarf companions in ten of thirteen systems. These systems provide a useful comparison to MWC 656 as they host rapidly rotating Be stars (–) with effective temperatures of 19–27 kK (spectral types B1–B3 with luminosity classes III–V), similar to MWC 656, accompanied by stripped subdwarfs.
@@
Wang et al. identified subdwarf companions through narrow cross-correlation peaks with 45 kK O-type star templates. Applying the same method to MWC 656, we detected no such peak, suggesting a different type of stripped companion. In fact, while the Wang et al. subdwarfs have inferred temperatures around 45 kK and luminosities –, our best-fit model for MWC 656 requires a much hotter (85 kK) and ten times more luminous stripped star.
Among the Wang et al. sample, only HD 55606 and HD 157832 show possible wind signatures and variability in N v, though weaker than in MWC 656 (see Fig. 9). HD 55606 was first classified as a Be+subdwarf binary based on double-peaked He i emission and weak He ii absorption in the optical (Chojnowski et al., 2018). He ii in absorption in HD 55606, versus emission in MWC 656, could be due to the higher inferred mass and temperature of the companion in MWC 656 (1.5 M⊙ compared to 0.92 M⊙). In contrast, HD 157832 shows N v variability but cross-correlation yielded no detection of a companion by Wang et al.. Notably, HD 157832 is a known X-ray source ( erg s-1; Lopes de Oliveira and Motch, 2011), classifying it as a -Cas analogue. X-ray emission in -Cas-like systems has been proposed to arise from stripped star winds interacting with the Be-star disc (Langer et al., 2020a), a scenario that may also apply to MWC 656.
@@
Figure 10 compares MWC 656 to known stripped stars in the Hertzsprung-Russell diagram. The MWC 656 companion is hotter and more luminous than the Wang et al. subdwarfs, and hotter but similarly luminous compared to the recently identified intermediate-mass stripped stars in the LMC and SMC (Drout et al., 2023). This aligns with expectations that lower metallicity environments (like the Magellanic Clouds) lead to less efficient envelope stripping and cooler, only partially stripped stars (Klencki et al., 2022; Dutta and Klencki, 2024). The derived mass-loss rate is in good agreement with the Vink (2017) prescription for stripped helium stars, as also found for partially stripped stars by Ramachandran et al. (2024), although we note that these studies focus on somewhat larger, cooler stars with lower surface gravities. In contrast, other detected stripped stars in the Magellanic Clouds are reported to have winds that are orders of magnitude weaker based on optical diagnostics (Götberg et al., 2023). We also include the helium-dominated merger products CD–46 8926 and CD–51 11879 (Krtička et al., 2024) in the figure, which are hot subdwarfs with kK and surface abundances pointing to a merger origin. Their luminosities and wind mass loss rates are an order of magnitude below those of MWC 656, yet they exhibit very similar UV wind features. In particular, CD–46 8926 shows a N v doublet closely resembling that of the stripped companion in MWC 656 (Krtička et al., 2024, their Fig. 5). For reference, we show the He-ZAMS for solar metallicity in Fig. 10, calculated by Picco et al. (2024) using MESA evolutionary models (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). The MWC 656 companion lies close to this theoretical boundary, although its inferred luminosity-to-mass ratio exceeds that expected for core-helium burning. The inferred luminosity implies an additional contribution, likely due to He-shell burning, suggesting that the companion may already have evolved beyond the core-He-burning phase. Detailed binary evolution models will be required to test this interpretation.

### v2

@@ 5 Signs of the companion in the UV
The HST/STIS FUV spectra, obtained at four orbital phases (two near opposite quadratures at , and two near conjunction, when the Be star was in front of the companion along the line of sight at ; see Fig. 2), are displayed in Fig. 4. The Be star model reproduces most of the FUV spectrum well, confirming that the observed flux is dominated by the Be star. However, several features are not reproduced by the model and instead suggest the presence of a hotter component in the system. We discuss these features in the following.
@@ 5.1 He ii emission
The He ii line, shown in the rightmost zoom panel of Fig. 4, displays clear phase-dependent variability (see also Fig. 11). Mid- to late-type B stars generally do not exhibit He ii lines, and therefore the variability seen here is unlikely to originate from the Be star. Despite the limited number of available epochs, we detect an emission component that shifts in anti-phase with the Be star, consistent with the companion (see Appendix B). This behaviour mirrors our findings in the optical, where the He ii line also appears in emission, tracking the companion.
@@ 5.2 High-ionisation wind lines
Beyond He ii, the spectra show several high-ionisation species, shown in Fig. 4. The N v 1239, 1243 doublet exhibits prominent P Cygni profiles, especially near quadrature, with blueshifted absorption and redshifted emission components. This suggests a hot, radiatively driven wind. Although the Si iv 1393, 1402 and C iv 1548, 1551 lines do not show clear emission components, they exhibit sloped blueshifted absorption troughs indicative of wind outflows, especially at and . These wind features coincide with an interstellar absorption component.
The Si iv and C iv lines can arise from super-ionisation in B-star winds, typically interpreted as the result of shock-generated X-rays that raise ionisation levels above those expected for the stellar temperature (e.g. Puls et al., 2008). Such effects are common in Be stars and explain their stronger UV wind lines compared to non-emission B stars (Grady et al., 1987; Prinja, 1989). Indeed, incorporating the observed level of X-rays into our 21 kK Be-star model reproduces the absorption components of the Si iv and C iv resonance lines. However, even accounting for super-ionisation, N v P-Cygni emission is only observed in the most luminous Be stars earlier than B1 (Kogure and Hirata, 1982; Rivinius et al., 2013). As discussed in Sect. 8.1, no comparable N v emission is seen in other Be stars of similar type to MWC 656. Producing such features in MWC 656 therefore requires either a substantially hotter ionising source or an unusually high X-ray luminosity.
@@ 5.3 Doppler shifts
Using the orbital solution presented in Sect. 3.2, we predicted Doppler shifts for both components across the observations. For the Be star, the expected maximum RV amplitude between observations is approximately 26 km s-1. To estimate RV shifts from the observed UV spectra, we performed an iterative cross-correlation of each epoch against the spectrum, focussing on wavelength regions free from strong interstellar absorption and wind lines. This yields relative RVs of , , and km s-1 for , and , respectively. The observed shift at is qualitatively consistent with the expected lower RV at this phase, although the measured amplitude is smaller than predicted. However, we note that, given the moderate S/N, broad spectral features, and relatively low RV amplitude of the Be star, precise measurements are challenging. The companion is predicted to have an RV amplitude of about 140 km s-1 between observations. However, the FUV spectrum is dominated by the Be star, making it challenging to directly measure the companion’s RVs. Nevertheless, we recover He ii emission from the companion through spectral disentangling using the predicted orbital velocities (Fig. 11), providing indirect support for the expected motion. A puzzling feature of the UV observations is the apparent Doppler shift seen in the high-ionisation wind lines, which follows neither of the binary components and is difficult to explain in any binary companion scenario. We address this aspect in more detail in Sect. 8.2.
@@ 5.4 Phase-dependent spectral variability
The four epochs of HST/STIS FUV spectroscopy reveal phase-dependent variability in the high-ionisation UV wind lines. Resonance line profiles appear to strengthen near quadrature, with, for example, the C iv doublet showing more prominent blueshifted absorption troughs during these phases, while features weaken near conjunction (–0.17). Variability in the line profiles of resonance lines such as Si iv and C iv is common in Be stars (e.g. Grady et al., 1987). However, the observed changes in the UV line profile may also reflect the interaction between the companion and the Be disc. In binaries, hydrodynamic simulations show that the companion can perturb the Be disc, exciting spiral density waves that produce phase-dependent variations in emission and absorption (Panoglou et al., 2016; Rubio et al., 2025). The phase-dependence could also indicate ionisation of the Be star wind by the hot companion, in which case the absorption would originate on the far side of the system during conjunction, or the wind of the hot companion itself, whose signatures diminish when it is partially obscured behind the Be star. However, we note that with only four UV epochs, some apparent phase-dependence could be coincidental.
@@ 6 Composite model with a stripped-star companion
The UV data revealed spectral features, such as high ionisation wind lines in N v and He ii emission, absent in the Be star spectrum, which point towards the presence of a hotter luminous companion in the system (Sect. 5). Motivated by these findings, we modelled the binary as a composite of two luminous stars, featuring the Be star and a stripped helium star companion. We performed a combined spectroscopic and photometric analysis, with a setup similar to that in Sect. 4, with the companion star parameters optimised to reproduce the observed UV wind lines.
@@
The companion’s effective temperature is primarily constrained by the observed UV emission lines. The presence of N v requires a minimum temperature of kK. However, reproducing the observed N v emission without accompanying strong C iv or Si iv emission requires significantly higher temperatures, kK, consistent with trends seen in Wolf-Rayet stars (e.g. Willis et al., 1986; Hamann and Gräfener, 2004; Hamann et al., 2006; Sander et al., 2025).
We explored companion temperatures from 65 to 95 kK, with corresponding scaled between 4.6 and 5.6, to be consistent with the helium zero-age main sequence (ZAMS). Models with lower temperatures produce stronger C iv and Si iv lines but weaker N v features. In contrast, models that are too hot have a steep SED, which reduces their contribution to the optical spectrum. By balancing the strength and shape of the N v P Cygni profile with the contribution to optical lines such as He ii 4686, our models suggest a best-fit temperature of kK. The sensitivity of the fit to the assumed temperature is illustrated in Appendix E, where we compare cooler and hotter companion models to the observations.
The resulting best-fit composite model is shown in Fig. 6, with key parameters summarised in Table 2. We derive mag and . We note that the corresponding -band extinction of mag agrees well with 3D dust map predictions (e.g. from Bayestar19; Green, 2018; Green et al., 2019), suggesting negligible system-intrinsic extinction. The Be star has an inferred luminosity of and a stellar radius of , the latter determined using the Stefan-Boltzmann law. As expected, attributing part of the observed flux to the stripped companion results in slightly lower inferred luminosity and radius compared to the single-star model (Sect. 4), but the differences are minimal. This is because the companion’s contribution to the total flux is small, and the Be star dominates both observed UV and optical light.
@@ 7.1 Companion mass
The mass of the companion in MWC 656 can be constrained by combining the dynamical mass ratio with spectroscopic and evolutionary estimates of the mass of the Be star.
@@
which is illustrated in Fig. 7. The Be star mass inferred from our composite spectroscopic model is (The mass obtained from the single-star analysis was 8 ). Combining this with the mass function implies an inclination of degrees and a companion mass of .
An evolutionary mass was also derived by comparing the effective temperature and luminosity of the Be star with simulated evolutionary tracks. We used the BONNSAI444The BONNSAI web-service is available at www.astro.uni-bonn.de/stars/bonnsai code (Schneider et al., 2014), which implements a Bayesian statistical method to determine probability distributions of fundamental stellar parameters. We used the massive star evolutionary model grids by Brott et al. (2011) and flat prior distributions in stellar mass and age. This yields , which corresponds to a companion mass of , consistent with the spectroscopic mass estimate. However, this value may represent an upper limit, as past mass accretion could make the Be star temporarily appear over-luminous relative to its true mass (e.g. Kippenhahn and Meyer-Hofmeister, 1977; Lau et al., 2024).
Together, dynamical, spectroscopic, and evolutionary constraints consistently place the companion mass between 1.0 and 1.9 M⊙. Two key conclusions follow: First, a BH companion is strongly disfavoured. The inferred mass lies well below the canonical lower mass threshold for BHs (3 M⊙). For comparison, C14 proposed a BH mass of 3.8–6.9 M⊙. Even accommodating the lowest of these values would require a Be star mass of 19 M⊙, which is ruled out at by both spectroscopic and evolutionary estimates. Second, a white dwarf companion is also unlikely. Even the minimum allowed dynamical companion mass at ( M⊙) implies a very massive white dwarf, and the most likely companion mass of M⊙ exceeds the Chandrasekhar limit (1.4 M⊙), placing most of the plausible mass range above this threshold.
@@ 7.2 Could the companion still be an accreting compact object?
A white dwarf accretor provides another possible explanation. X-rays from accretion could photo-ionise the Be-star wind and produce N v emission. This scenario has been proposed to explain the so-called Cas stars, Be stars with anomalously high X-ray luminosities that may host accreting white dwarfs (e.g. Smith et al., 2016; Gies et al., 2023). Although MWC 656 is significantly less X-ray luminous than typical Cas stars, we tested this scenario by incorporating increasing levels of X-ray heating in the Be-star model (Sect. 5.2). Reproducing the observed N v strength requires X-ray luminosities about two orders of magnitude higher than observed in MWC 656, and the models simultaneously predict strong C iv and Si iv emission that is not seen in the data.
@@ 8.1 Comparison of MWC 656 to known stripped stars
To place our results in context, we compare the inferred properties of the stripped star in MWC 656 to those of known stripped stars. These hot, compact helium stars are formed by envelope stripping in binary systems. Stripped stars span a wide mass range, from low-mass subdwarf B and O stars (; e.g. Heber, 2009; Mourard et al., 2015; Klement et al., 2022) to massive Wolf-Rayet stars (; e.g. Shenar et al., 2016, 2018, 2020b; Sander and Vink, 2020), with several intermediate-mass cases recently identified in the Magellanic Clouds (Ludwig et al., 2025; Drout et al., 2023; Ramachandran et al., 2023, 2024).
Stripped stars are often found with rapidly rotating main-sequence companions, believed to have spun up via accretion during envelope stripping. Wang et al. (2021), for example, identified Galactic Be+subdwarf candidates using HST UV spectroscopy, detecting hot subdwarf companions in ten of 13 systems. These systems provide a useful comparison to MWC 656 as they host rapidly rotating Be stars (–) with effective temperatures of 19–27 kK (spectral types B1–B3 with luminosity classes III–V), similar to MWC 656, accompanied by stripped subdwarfs.
@@
Wang et al. identified subdwarf companions through narrow cross-correlation peaks with 45 kK O-type star templates. Applying the same method to MWC 656, we did not detect such a peak, suggesting a different type of stripped companion. In fact, while the Wang et al. subdwarfs have inferred temperatures around 45 kK and luminosities –, our best-fit model for MWC 656 requires a much hotter (85 kK) and ten times more luminous stripped star.
Among the Wang et al. sample, only HD 55606 and HD 157832 show possible wind signatures and variability in N v, though weaker than in MWC 656 (see Fig. 9). HD 55606 was first classified as a Be+subdwarf binary based on double-peaked He i emission and weak He ii absorption in the optical (Chojnowski et al., 2018). He ii in absorption in HD 55606, versus emission in MWC 656, could be due to the higher inferred mass and temperature of the companion in MWC 656 (1.5 M⊙ compared to 0.92 M⊙). In contrast, HD 157832 shows N v variability but cross-correlation did not yield a companion detection by Wang et al.. Notably, HD 157832 is a known X-ray source ( erg s-1; Lopes de Oliveira and Motch, 2011), classifying it as a -Cas analogue. X-ray emission in -Cas-like systems has been proposed to arise from stripped star winds interacting with the Be-star disc (Langer et al., 2020a), a scenario that may also apply to MWC 656.
@@
Figure 10 compares MWC 656 to known stripped stars in the Hertzsprung-Russell diagram. The MWC 656 companion is hotter and more luminous than the Wang et al. subdwarfs, and hotter but similarly luminous compared to recently identified intermediate-mass stripped stars in the Large and Small Magellanic Clouds (Drout et al., 2023). This aligns with expectations that lower-metallicity environments (like the Magellanic Clouds) lead to less efficient envelope stripping and cooler, only partially stripped stars (Klencki et al., 2022; Dutta and Klencki, 2024). The derived mass-loss rate is in good agreement with the Vink (2017) prescription for stripped helium stars, as also found for partially stripped stars by Ramachandran et al. (2024), although we note that these studies focus on somewhat larger, cooler stars with lower surface gravities. In contrast, other detected stripped stars in the Magellanic Clouds are reported to have winds that are orders of magnitude weaker based on optical diagnostics (Götberg et al., 2023). We also include the helium-dominated merger products CD–46 8926 and CD–51 11879 (Krtička et al., 2024) in the figure, which are hot subdwarfs with kK and surface abundances pointing to a merger origin. Their luminosities and wind mass loss rates are an order of magnitude below those of MWC 656, yet they exhibit very similar UV wind features. In particular, CD–46 8926 shows a N v doublet closely resembling that of the stripped companion in MWC 656 (Krtička et al., 2024, their Fig. 5). For reference, we show the helium ZAMS for solar metallicity in Fig. 10, calculated by Picco et al. (2024) using MESA evolutionary models (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). The MWC 656 companion lies close to this theoretical boundary, although its inferred luminosity-to-mass ratio exceeds that expected for core-helium burning. The inferred luminosity implies an additional contribution, likely due to He-shell burning, suggesting that the companion may already have evolved beyond the core-He-burning phase. Detailed binary evolution models will be required to test this interpretation.